Consider the language L _ {w E {a,b) : the longest run of as in u, is longer than any run of bs in w} For example, abbbbaabbbaaaaa E L because the longest run of bs in it has length four, while the longest run of as has length six. Prove that L is not context-free.
Answer #1

To Prove a language is not Context Free we use Pumping Lemma for CFL.

If L is any CFL and z \small \epsilon L such that |z| \tiny \geq n, if z = uvwxy \small \epsilon L ( |uvx| \tiny \leq |z| , |vx| != 0)

then uviwxiy \small \epsilon L , \small \forall i \tiny \geq 0.

Every CFL satisfies pumping lemma.

Now consider z = abbbbaabbbaaaaaa

take u = a, v = bbbb, w = aa, x = bbb, y = aaaaaa

Here in uviwxiy take i = 2 now it will bw uv2wx2y

Hence uv2wx2y = a(bbbb)2aa(bbb)2aaaaaa = abbbbbbbbaabbbbbbaaaaaa

Now in this string it is saying that longest run of a's in w is shorter than any run of b's in w.

This is a contradiction, Hence this Language is not CFL.